# Rich Semantic Knowledge Enhanced Large Language Models for Few-shot Chinese Spell Checking

### v1

@@ 5.4 Experimental Results
Table 3 shows the evaluation results of our RS-LLM against other models on the SIGHAN15, LAW, MED and ODW test sets. We found that the experimental results of LLM have certain fluctuations with temperature changes, so our baseline experiments of LLM and RS-LLM experiments on the same model all use the same temperature. It is observed that our RS-LLM outperforms all baselines on all metrics in most cases for all datasets. In particular, LLM uses the same few-sample cues as RS-LLM but does not add semantic information when performing the CSC task. The experimental results of this paper verify the effectiveness and superiority of LLM semantic information in CSC tasks.
As shown in the table reftable:result, our spell checking method, RS-LLM over glm-3-turbo, achieves 0.6% error detection and 2.1% correction on SIGHAN15 dataset using Rich Semantic compared to standard LLM. And RS-LLM on glm-3-turbo registers 1% error detection and 2.4% correction rate were achieved on the MED dataset. RS-LLM on gpt-3.5-turbo, achieved a 1.4% correction rate of 1.4% error detection on SIGHAN 15 using Rich Semantic. In the LAW dataset, RS-LLM on gpt-3.5-turbo beats gpt-3.5-turbo by 2.8% in detection and 2.6% in correction, while RS-LLM on gpt-3.5-turbo outperforms gpt-3.5-turbo by 3.1% and 1.8% in the MED dataset. On the ODW dataset, our method also shows significant gains, with RS-LLM on gpt-3.5-turbo leading gpt-3.5-turbo by 4.9% in detection and 5.8% in correction.
In addition, it’s clear that the CSC model based on BERT performs worse than LLM in few-shot scenarios, both at the detection level and the correction level. The ability of the CSC model based on BERT to detect and correct incorrect characters excessively relies on the training data of CSC, especially the ability to correct incorrect characters.
@@ 5.5.1 Impact of Different In-context Learning Approaches
Fig.4 and Fig.5 show the evaluation results of our RS-LLM and LLM with different context learning methods. For specific experimental results, see the table 4 in the appendix 5. From Fig.4 and Fig.5, we find that the performance of the CSC task on LLM improves on most datasets as the number of examples in the context increases, both in terms of detection and correction. Furthermore, RS-LLM is more effective than LLM in few-shot scenarios compared to zero-shot and one-shot scenarios. Such experimental results reflect the effectiveness of our in-context learning approach designed for CSC task.
5.5.2 Impact of Different Semantic Information
In order to further study the impact of semantic information on LLMs when performing CSC tasks. We design to use phonetic information, radical information, structural information and strokes information. Each of these four types of information is added to our zero-shot prompt, one-shot prompt, and few-shot prompt. To explore the impact of these four prompts on the test set of SIGHAN15, LAW, MED, and ODW in three different contexts of learning. The specific experimental results is Table 5 in Appendix 5.
Experimental results show that single speech, radical, structure and stroke information can improve CSC task on most datasets in few-shot scenarios. It is worth noting that speech and radical information are basically enhanced on all datasets, and structural and stroke information can improve the performance of LLM on the CSC task on some datasets. But this improvement is relatively small compared to other information.
@@ 6 Conclusion and Future Work
In this paper, we introduce an in-context learning method named RS-LLM for CSC task, one of whose core components is using Chinese Rich Semantics in LLMs for CSC task. RS-LLM utilizes adding a small number of specific Chinese rich semantic information into a specific few-shot prompt set for CSC task, aiming to allow LLMs to have a fuller understanding of the semantics when doing CSC task. Experimental results show that this LLMs-independent approach can help existing LLMs better recognize and correct phonetically and visually erroneous characters in CSC tasks. Considering the impact of the similarity of errors in the few-shot prompt and errors in the sentence on LLMs’ understanding of the sentence when LLMs perform CSC task. Next we will try to enhance our in-context learning method through similarity retrieval when constructing few-shot prompts.
@@
To verify the effectiveness of RS-LLM, we conducted extensive experiments on two benchmark datasets of different domains and scales. The results indicate that RS-LLM delivers SOTA results in few -shot scenarios. However, since most of the errors in the CSC dataset are attributed to visual and phonetical errors, we incorporate phonetic and radical information into the prompt template. However, it is difficult to ensure that the existing manually formulated prompt templates are optimal, and the optimal prompt sentences for CSC require further research. Furthermore, relevant examples have been carefully selected to enable LLMs to identify potential visual and speech errors in a small number of scenarios. No consideration was given to the ability to motivate the LLMs through semantic similarity. We recognize these two limitations and plan to address them in future research efforts.
@@ Appendix A The Detailed Experimental Results
zero-shot SIGHAN15 glm-3-turbo 47.1 28.7 29.4 29.0 44.1 22.7 23.3 23.0
RS-glm-3-turbo 47.9 26.5 27.7 27.1 45.8 22.4 23.5 22.9
gpt-3.5-turbo 33.6 22.7 36.6 28.1 31.8 20.4 32.9 25.2
RS-gpt-3.5-turbo 42.5 25.8 34.9 29.7 40.4 22.8 30.9 26.2
LAW glm-3-turbo 36.4 20.7 26.7 23.3 34.4 17.6 22.8 19.9
RS-glm-3-turbo 60.8 42.7 40.4 41.5 57.6 36.1 34.1 35.1
gpt-3.5-turbo 28.4 20.6 34.1 25.7 26.4 18.2 30.2 22.7
RS-gpt-3.5-turbo 44.8 26.4 32.2 29.0 42.6 22.8 27.8 25.1
MED glm-3-turbo 49.0 26.9 35.4 30.6 46.6 22.9 30.1 26.0
RS-glm-3-turbo 53.4 26.5 30.9 28.6 51.2 22.4 26.1 24.1
gpt-3.5-turbo 33.8 19.6 34.9 25.1 31.6 16.9 30.1 21.6
RS-gpt-3.5-turbo 36.6 17.8 28.3 21.9 33.6 13.7 21.7 16.8
ODW glm-3-turbo 54.8 45.8 30.9 36.9 52.4 39.0 26.3 31.4
RS-glm-3-turbo 69.6 58.1 45.0 50.8 66.6 50.7 39.3 44.3
gpt-3.5-turbo 38.8 31.3 50.7 38.7 35.6 27.5 44.7 34.1
RS-gpt-3.5-turbo 57.2 41.2 43.5 42.3 53.6 34.7 36.6 35.6
one-shot SIGHAN15 glm-3-turbo 43.7 28.4 40.1 33.3 39.8 22.8 32.2 26.7
RS-glm-3-turbo 42.8 23.6 29.8 26.3 40.6 20.0 25.1 22.2
gpt-3.5-turbo 29.6 21.2 36.6 26.9 27.7 18.9 32.7 24.0
RS-gpt-3.5-turbo 38.6 23.2 33.1 27.3 36.6 20.4 29.2 24.1
LAW glm-3-turbo 61.6 47.7 60.0 53.1 57.2 40.8 51.4 45.5
RS-glm-3-turbo 60.8 43.1 48.6 45.7 57.8 37.9 42.8 40.2
gpt-3.5-turbo 34.6 24.8 40.4 30.8 32.2 21.9 35.7 27.2
RS-gpt-3.5-turbo 41.6 24.3 32.2 27.7 39.6 21.3 28.2 24.3
MED glm-3-turbo 52.6 31.8 46.9 37.9 50.0 27.9 41.2 33.3
RS-glm-3-turbo 52.8 29.7 39.4 33.8 49.2 23.7 31.4 27.0
gpt-3.5-turbo 16.0 11.8 24.8 16.0 14.2 9.9 20.8 13.5
RS-agpt-3.5-turbo 25.4 13.8 25.2 17.8 24.0 12.1 22.1 15.7
ODW glm-3-turbo 80.0 69.5 68.7 69.1 76.4 62.6 61.8 62.2
RS-glm-3-turbo 74.2 62.5 54.6 58.3 70.2 53.7 47.0 50.1
gpt-3.5-turbo 40.8 31.9 50.0 38.9 38.6 29.2 45.8 35.7
RS-gpt-3.5-turbo 68.2 53.4 48.1 50.6 65.6 47.9 43.1 45.4
few-shot SIGHAN15 glm-3-turbo 45.6 28.9 40.1 33.6 42.2 23.9 33.1 27.7
RS-glm-3-turbo 46.4 29.7 40.1 34.2 43.5 25.3 34.2 29.1
gpt-3.5-turbo 37.0 24.1 38.6 29.7 35 21.6 34.6 26.6
RS-gpt-3.5-turbo 46.0 27.3 36.0 31.1 44.6 25.1 33.1 28.5
LAW glm-3-turbo 64.4 56.8 49.0 52.6 62.0 51.4 44.3 47.6
RS-glm-3-turbo 64.6 56.3 49.0 52.4 61.8 50.5 43.9 47.0
gpt-3.5-turbo 46.6 31.9 45.9 37.6 43.8 28.1 40.4 33.1
RS-gpt-3.5-turbo 56.2 37.6 43.9 40.5 54.0 33.9 39.6 36.5
MED glm-3-turbo 57.0 34.7 49.6 40.8 52.6 27.9 39.8 32.8
RS-glm-3-turbo 59.4 37.4 47.4 41.8 56.0 31.5 39.8 35.2
gpt-3.5-turbo 30.8 17.5 32.3 22.7 28.8 15.1 27.9 19.6
RS-gpt-3.5-turbo 39.0 20.7 34.1 25.8 36.4 17.2 28.3 21.4
ODW glm-3-turbo 81.0 69.9 69.1 69.5 76.6 61.4 60.7 61.0
RS-glm-3-turbo 78.0 67.2 64.1 65.6 73.4 58.0 55.3 56.6
gpt-3.5-turbo 65.0 52.3 66.0 58.0 61.8 47.4 59.9 52.9
RS-gpt-3.5-turbo 75.6 63.7 62.2 62.9 73.4 58.4 58.0 58.7
Table 4: The performance of different In-context Learning approaches added in prompts on LLMs. RS-gpt-3.5-turbo means RS-LLM on gpt-3.5-turbo and RS- glm-3-turbo means RS-LLM on glm-3-turbo.
Dataset Method Detection Level Correction Level
Acc. Pre. Rec. F1 Acc. Pre. Rec. F1
SIGHAN15 gpt-3.5-turbo 37.0 24.1 38.6 29.7 35.0 21.6 34.6 26.6
gpt-3.5-turbo+拼音 43.8 26.3 35.5 30.2 42.2 23.8 32.2 27.4
gpt-3.5-turbo+部首 44.3 25.7 34.2 29.3 41.5 22.9 30.5 26.2
gpt-3.5-turbo+结构 40.1 20.7 26.1 23.1 38.6 18.2 22.9 20.3
gpt-3.5-turbo+笔画 43.0 24.7 32.0 27.9 40.6 20.8 27.0 23.5
@@
LAW gpt-3.5-turbo 46.6 31.9 45.9 37.6 43.8 28.1 40.4 33.1
gpt-3.5-turbo+拼音 54.6 35.3 42.8 38.7 52.6 32.0 38.8 35.1
gpt-3.5-turbo+部首 53.8 34.9 42.8 38.5 52.0 32.1 39.2 35.3
gpt-3.5-turbo+结构 54.0 35.8 43.9 39.4 52.0 32.6 40.0 35.9
gpt-3.5-turbo+笔画 51.4 34.3 45.5 39.1 48.8 30.5 40.4 34.7
RS-gpt-3.5-turbo 56.2 37.6 43.9 40.5 54.0 33.9 39.6 36.5
MED gpt-3.5-turbo 30.8 17.5 32.3 22.7 28.8 15.1 27.9 19.6
gpt-3.5-turbo+拼音 35.4 18.5 31.4 23.3 33.4 15.9 27.0 20.0
gpt-3.5-turbo+部首 34.0 17.7 30.5 22.4 32.0 15.1 26.1 19.2
gpt-3.5-turbo+结构 36.6 19.6 33.2 24.7 34.4 16.8 28.3 21.1
gpt-3.5-turbo+笔画 23.0 12.8 24.3 16.7 21.6 11.1 21.2 14.6
RS-gpt-3.5-turbo 39.0 20.7 34.1 25.8 36.4 17.2 28.3 21.4
ODW gpt-3.5-turbo 65.0 52.3 66.0 58.0 61.8 47.4 59.9 52.9
gpt-3.5-turbo+拼音 71.8 57.5 57.3 57.3 69.2 52.5 52.3 52.4
gpt-3.5-turbo+部首 73.0 59.5 58.4 58.9 70.4 54.5 53.4 53.9
gpt-3.5-turbo+结构 71.0 55.6 59.2 59.2 67.8 49.8 53.1 51.4
gpt-3.5-turbo+笔画 73.4 59.2 60.3 59.7 70.4 53.6 54.6 54.1
RS-gpt-3.5-turbo 75.6 63.7 62.2 62.9 73.4 58.4 58.0 58.7
Table 5: The performance of different semantic information is added to few-shot prompts on LLMs. ’拼音’ means phonetic information, ’部首’ means radical information, ’结构’ means structural information, ’笔画’ means strokes information.

### v3

@@ 5.4 Experimental Results
Table 3 shows the evaluation results of our RS-LLMs comparing to other models on the test sets of SIGHAN15, LAW, MED, and ODW. It is observed that our RS-LLM consistently outperforms all baselines on all datasets in all metrics. Especially compared with LLM, LLM uses the same few-shot prompt as RS-LLM but does not add semantic information when performing CSC tasks. The experimental results of this period verify the effectiveness and superiority of our semantic information on LLM for CSC task.
From time to time experiments, we found that the experimental results are not completely consistent because the performance of online APIs varies. Therefore, we use $\pm$ in the results to explain the error band of the experiment.
As shown in Table 3, our spell-checking approach, RS-LLM on ChatGLM2-6B, achieves a notable 8.0% improvement in error detection and an 9.8% increase in correction on SIGHAN 15 using Rich Semantic, compared to the standard LLM. Impressively, RS-LLM on gpt-3.5-turbo registers a 9.9% boost in F1-score for detection and a 12.3% leap in correction. In the LAW dataset, RS-LLM on ChatGLM2-6B beats ChatGLM2-6B by 5.8% in detection and 6.6% in correction, while RS-LLM on gpt-3.5-turbo outperforms gpt-3.5-turbo by 16.9% and 14.6%, respectively. The trend continues in the MED dataset, where RS-LLM on ChatGLM2-6B surpasses ChatGLM2-6B by 13.1% in detection and 9.3% in correction, and RS-LLM on gpt-3.5-turbo exceeds gpt-3.5-turbo by 15.3% and 13.7%. On the ODW dataset, our method also shows significant gains, with RS-LLM on ChatGLM2-6B leading ChatGLM2-6B by 6.9% in detection and 5.6% in correction, and RS-LLM on gpt-3.5-turbo outdoing gpt-3.5-turbo by 14.1% and 12.7%.
@@ 5.5 Analyses and Discussions
In order to verify the effectiveness of the prompt sentence, we count the number of sentences whose length remained the same with and without a specific prompt, and the experimental results are shown in Table 2. It is obvious that LLM performs better than most of the BERT-based CSC models at both the detection and correction levels. The ability of the BERT-based CSC model to detect and correct erroneous characters is overly dependent on the training data of CSC, especially the ability to correct erroneous characters. The LLMs based approach shows better generalization in the few-shot scenario. However, we find that LLM’s performance on the CSC task heavily depends on the foundation model. When the scale of parameters of the foundation model is larger, the model performs better on the CSC task. Although the performance of RS-LLM cannot currently outperform Scope, we believe that with the continuous update of the base model performance, the RS-LLM method will continue to improve and show a better improvement trend.
@@ 5.5.1 Impact of Different In-context Learning Approaches
Fig.4 and Fig.5 show the evaluation results of different in-context learning approaches. From Fig.4 and Fig.5, we discover that the performance of CSC task on LLMs improves as the number of examples increasing, both in terms of detection and correction. Additionally, it’s clear that RS-LLM is more effective than most BERT-based models in terms of zero-shot, one-shot, and few-shot scene. Such experimental results reflect the effectiveness of our in-context learning strategy designed for CSC task. The specific experimental results are shown in Table 5 in Appendix 5.
5.5.2 Impact of Different Chinese Rich Semantic Information
In order to further study the impact of semantic information on LLMs when performing CSC tasks. We design to use phonetic information, radical information, structural information and strokes information. Each of these four types of information is added to our zero-shot prompt, one-shot prompt, and few-shot prompt. To explore the impact of these four prompts on the test set of SIGHAN15, LAW, MED, and ODW. The specific experimental results are Table 4 in Appendix 5.
The experimental results show that individual phonetic, radical, structural, and stroke information leads to improvements in CSC task across all datasets. Notably, the phonetic and radical information contribute the most significant enhancements, followed by structural information. While strokes information does show some improvement, it’s relatively puny compared to the others.
@@ 6 Conclusion and Future Work
In this paper, we introduce an in-context learning method named RS-LLM for CSC task, one of whose core components is using Chinese Rich Semantics in LLMs for CSC task. RS-LLM utilizes adding a small number of specific Chinese rich semantic information into a specific few-shot prompt set for CSC task, aiming to allow LLMs to have a fuller understanding of the semantics when doing CSC task. Experimental results show that this LLMs-independent approach can help existing LLMs better recognize and correct phonetically and visually erroneous characters in CSC tasks. Considering the impact of the similarity of errors in the few-shot prompt and errors in the sentence on LLMs’ understanding of the sentence when LLMs perform CSC task. In the future, we will try to construct dynamic prompt for each sample through semantic similarity retrieval.
@@
To verify the effectiveness of RS-LLM, we conducted extensive experiments on two benchmark datasets of different domains and scales. The results indicate that RS-LLM delivers SOTA results in few-shot scenarios. Since most of the errors in the CSC dataset are attributed to visual and phonetical errors, we incorporate phonetic and radical information into the prompt template. However, it is difficult to ensure that the existing manually formulated prompt templates are optimal, and the optimal prompt sentences for CSC require further research. Furthermore, relevant examples have been carefully selected to enable LLMs to identify potential visual and speech errors in a small number of scenarios. No consideration was given to the ability to motivate the LLMs through semantic similarity. We recognize these two limitations and plan to address them in future research efforts.
@@ Acknowledgments
This work was partly supported by China Postdoctoral Science Foundation (No. 2023M731253) and Hubei Provincial Natural Science Foundation (No. 2023AFB487).
@@ Appendix A The Detailed Experimental Results
Dataset Method Detection Level Correction Level
Acc. Pre. Rec. F1 Acc. Pre. Rec. F1
SIGHAN15 gpt-3.5-turbo 36.4 22.3 33.0 26.6 34.3 19.3 28.6 23.1
gpt-3.5-turbo+拼音 45.7 31.2 46.7 37.4 43.0 27.5 41.2 33.0
gpt-3.5-turbo+部首 48.1 29.7 36.8 32.8 45.6 26.6 31.8 28.4
gpt-3.5-turbo+结构 45.2 27.8 39.4 32.6 43.6 25.4 36.0 29.8
gpt-3.5-turbo+笔画 38.7 23.3 35.3 28.1 36.5 20.4 30.9 24.5
RS-gpt-3.5-turbo 50.6 32.5 41.6 36.5 48.1 31.2 40.8 35.4
LAW gpt-3.5-turbo 48.8 30.2 38.8 34.0 46.2 26.2 33.7 29.5
gpt-3.5-turbo+拼音 64.0 45.4 48.6 47.0 60.8 39.6 42.4 40.9
gpt-3.5-turbo+部首 65.0 45.7 49.4 47.5 61.4 39.1 42.4 40.7
gpt-3.5-turbo+结构 62.4 41.8 43.1 42.5 59.6 36.5 37.7 37.1
gpt-3.5-turbo+笔画 60.2 40.9 47.8 44.1 55.4 32.9 38.4 35.4
RS-gpt-3.5-turbo 64.6 46.6 56.1 50.9 60.8 40.4 48.6 44.1
MED gpt-3.5-turbo 41.4 16.8 30.1 21.5 38.6 13.3 23.9 17.1
gpt-3.5-turbo+拼音 52.6 31.0 37.9 34.1 48.2 20.9 32.5 25.5
gpt-3.5-turbo+部首 50.0 30.3 37.3 33.4 46.2 19.1 29.3 23.1
gpt-3.5-turbo+结构 46.6 26.0 31.5 28.5 44.1 16.6 26.9 20.5
gpt-3.5-turbo+笔画 42.4 24.9 30.5 27.5 40.0 16.5 28.1 20.8
RS-gpt-3.5-turbo 56.0 31.0 45.2 36.8 43.3 25.3 39.3 30.8
ODW gpt-3.5-turbo 63.0 45.8 50.0 47.8 59.2 39.2 42.8 40.9
gpt-3.5-turbo+拼音 70.4 55.7 58.0 56.8 67.2 49.8 51.9 50.8
gpt-3.5-turbo+部首 70.8 55.0 60.3 57.6 66.0 46.7 51.2 48.8
gpt-3.5-turbo+结构 70.8 55.2 58.4 56.8 66.2 46.9 49.6 48.2
gpt-3.5-turbo+笔画 69.4 53.4 54.6 54.0 65.8 46.6 47.7 47.2
RS-gpt-3.5-turbo 72.4 59.1 64.9 61.9 70.2 50.2 57.6 53.6
Table 4: The detailed impact of different rich semantic structures on few-shot learning. RS-gpt-3.5-turbo means RS-LLM on gpt-3.5-turbo. ’拼音’ means phonetic information, ’部首’ means radical information, ’结构’ means structural information, and ’笔画’ means strokes information.
@@
zero-shot SIGHAN15 gpt-3.5-turbo 25.9 15.6 26.1 19.6 23.9 13.2 22.0 16.1
RS-gpt-3.5-turbo 38.6 25.4 40.1 30.9 35.6 21.4 34.2 26.4
LAW gpt-3.5-turbo 32.2 19.9 30.6 24.1 30.2 17.4 26.7 21.0
RS-gpt-3.5-turbo 54.8 36.7 45.1 40.5 50.4 29.7 36.5 32.8
MED gpt-3.5-turbo 24.6 12.1 22.1 15.7 22.6 9.7 17.7 12.5
RS-gpt-3.5-turbo 43.0 25.1 41.6 31.3 40.0 21.1 35.0 26.3
ODW gpt-3.5-turbo 33.8 22.9 34.4 27.5 31.8 20.4 30.5 24.4
RS-gpt-3.5-turbo 60.6 50.0 51.3 50.6 55.7 41.1 42.1 41.6
one-shot SIGHAN15 gpt-3.5-turbo 34.5 19.2 29.0 23.1 32.9 17.1 25.9 20.6
@@
LAW gpt-3.5-turbo 41.2 26.1 37.3 30.7 38.6 22.5 32.2 26.5
RS-gpt-3.5-turbo 59.8 38.1 52.9 44.3 56.8 32.2 44.7 37.4
MED gpt-3.5-turbo 36.2 14.1 26.1 18.3 23.0 10.3 19.0 13.3
RS-gpt-3.5-turbo 51.4 30.0 44.3 35.8 47.2 23.7 35.0 28.3
ODW gpt-3.5-turbo 44.9 38.8 46.0 42.1 42.1 28.2 36.0 31.6
RS-gpt-3.5-turbo 64.8 49.5 58.8 53.8 61.8 44.7 53.1 48.5
few-shot SIGHAN15 gpt-3.5-turbo 36.4 22.3 33.0 26.6 34.3 19.3 28.6 23.1
RS-gpt-3.5-turbo 50.6 32.5 41.6 36.5 48.1 31.2 40.8 35.4
LAW gpt-3.5-turbo 48.8 30.2 38.8 34.0 46.2 26.2 33.7 29.5
RS-gpt-3.5-turbo 64.6 46.6 56.1 50.9 60.8 40.4 48.6 44.1
MED gpt-3.5-turbo 41.4 16.8 30.1 21.5 38.6 13.3 23.9 17.1
RS-gpt-3.5-turbo 56.0 31.0 45.2 36.8 43.3 25.3 39.3 30.8
ODW gpt-3.5-turbo 63.0 45.8 50.0 47.8 59.2 39.2 42.8 40.9
RS-gpt-3.5-turbo 72.4 59.1 64.9 61.9 70.2 50.2 57.6 53.6
Table 5: The detailed performance of different number of prompt samples. RS-gpt-3.5-turbo means RS-LLM on gpt-3.5-turbo.